# MCPVerse: An Expansive, Real-World Benchmark for Agentic Tool Use

### v1

@@ Introduction
  • Standard Mode: This mode is designed for the 64k-token context length, a standard supported by most state-of-the-art LLMs. We curated a set of $32$ MCPs (totaling $218$ tools), whose combined definitions occupy approximately 44k tokens. This reserves the remaining  20k tokens for the user prompt and the model’s response.
@@
  • Max-Scale Mode: This mode loads all $65$ MCPs with $552$ tools simultaneously, which collectively occupy a context length of approximately 140k tokens. In our evaluation, only Claude-4-Sonnet and Gemini-2.5-Pro were able to complete the evaluations in this setting due to its extensive context requirements.
Model Context Tools Oracle Mode Standard Mode Max-Scale Mode
Avg. L1 L2 L3 Avg. L1 L2 L3 Avg. L1 L2 L3
DeepSeek-V3-0324 $64\mathrm{k}$ - $49.55$ $66.66$ $46.51$ $35.48$ $37.58$ $58.33$ $30.23$ $24.19$ {-} {-} {-} {-}
DeepSeek-R1-0528 $64\mathrm{k}$ - $50.77$ $61.67$ $51.94$ $38.71$ $50.28$ $58.33$ $45.74$ $46.77$ {-} {-} {-} {-}
Claude-4-Sonnet $200\mathrm{k}$ - $\mathbf{57.81}$ $68.33$ $53.49$ $\mathbf{51.61}$ $\mathbf{61.01}$ $\mathbf{70.00}$ $\mathbf{56.59}$ $\mathbf{56.45}$ $\mathbf{57.77}$ $\mathbf{66.66}$ $\mathbf{55.04}$ $\mathbf{51.61}$
Qwen3-235B-A22B $128\mathrm{k}$ - $47.90$ $63.33$ $46.51$ $33.87$ $42.26$ $58.33$ $42.64$ $25.81$ {-} {-} {-} {-}
GPT-4o-20241120 $\mathrm{k} 128128$ $51.32$ $\mathbf{72.41}$ $46.46$ $35.09$ $26.83^{*}$ $53.33^{*}$ $5.50^{*} 11$$1.67^{*}$ {-} {-} {-} {-}
Qwen3-30B-A3B $128\mathrm{k}$ - $33.31$ $53.33$ $24.03$ $22.58$ $22.78$ $36.66$ $13.95$ $17.74$ {-} {-} {-} {-}
Gemini-2.5-Pro $1\mathrm{M}$ $512$ $48.25$ $65.00$ $42.64$ $37.10$ $47.33$ $61.67$ $48.06$ $32.26$ $42.28^{*}$ $60.00^{*}$ $42.64^{*}$ $24.19^{*}$
Kimi-K2-0711 $\mathrm{k} 128128$ $55.84$ $63.33$ $\mathbf{55.81}$ $48.39$ $16.34^{*}$ $25.00^{*}$ $24.03^{*}$ $0.00^{*}$ {-} {-} {-} {-}
Table 2: Accuracy results for all evaluated models across the Oracle, Standard, and Max-Scale modes. The table also lists each model’s context length (Context) and its native tool limit (Tools). Dashes (-) indicate that a model was not evaluated in a given mode due to its context length. Scores marked with were obtained using prompt-based function calling instead of native tool use due to context length limits.
The second challenge arises from tool number limits imposed by certain APIs. For example, the native tools parameter in the GPT-4o-20241120 and Kimi-K2-0711 API is capped at $128$ tools, which is insufficient for the $218$ tools in our Standard mode. The constraint is also present at a larger scale: the Gemini-2.5-Pro API’s limit of $512$ tools is inadequate for the $552$ tools in our Max-Scale Mode. To address these limitations, we employ a method we term prompt-based function calling. This approach involves providing the complete tool definitions directly within a custom system prompt (see details in Appendix 1), rather than using the native tools parameter, thereby bypassing the API’s fixed tool count limits.
Listing 1 Evaluation Pipeline
1toolset = []
2for mcp in mcp_list:
3 mc = connect_to(mcp)
4 toolset.append(mc.get_tools())
5
6task = get_user_input()
7history = [task]
8
9while True:
10 resp = LLM(history, toolset)
11 history.append(resp)
12 if not resp.tool_calls:
13 return
14
15 for tool_call in resp.tool_calls:
16 result = call_tool(tool_call)
17 history.append(result)
@@
Finally, we deliberately employ a simple evaluation pipeline. Our goal is to assess the fundamental capabilities of the LLM itself, rather than the effectiveness of a complex agentic framework like ReAct (Yao et al. 2023) or RolePlaying (Hu et al. 2025). These advanced methods often rely on sophisticated prompt engineering, which can obscure the model’s intrinsic abilities. Our direct approach, detailed in Listing 1, minimizes the influence of such artifacts and provides a clearer measure of the model’s core function calling proficiency. The hyperparameter settings for our experiments are detailed in the Appendix 4.
Results Analysis
@@
This section presents a detailed analysis of the experimental results. Owing to tool-mounting constraints, the Standard mode score for GPT-4o-20241120 and the Max-Scale mode score for Gemini-2.5-Pro are both obtained via prompt-based function calling.
Overall Model Performance. Based on the results presented in Table 2, Claude-4-Sonnet demonstrates definitive leadership in the Standard mode, achieving the highest average score of $61.01$. It consistently outperforms the other models across all three complexity levels, with its advantage most pronounced on L3 tasks, where it exceeds the non–prompt-based cohort’s average by $27.10$. DeepSeek‑R1‑0528 secures a solid second place with an average score of $50.28$, closely followed by Gemini‑2.5‑Pro at $47.33$. Following this, DeepSeek-V3-0324 ($37.58$) and Qwen3-235B-A22B ($42.46$) exhibit moderate performance, positioning them in the middle tier. At the lower end of the performance spectrum are Qwen3‑30B‑A3B ($22.78$), GPT‑4o ($26.83$), and Kimi‑K2‑0711 ($16.34$). The performances of GPT‑4o and Kimi‑K2‑0711 appear to be significantly hindered by their reliance on prompt‑based function calling due to API limitations, which likely conflicts with their native tool‑use training methodologies. Meanwhile, the performance of Qwen3‑30B‑A3B is likely constrained by its smaller model size, which limits its capacity for sophisticated tool analysis.
In Oracle mode, the performance gap between models narrows. Claude-4-Sonnet maintains a lead, achieving an overall score of $57.81$, closely followed by Kimi at $55.84$. Following them, a large cohort of models—including GPT-4o-20241120, DeepSeek-R1-0528, DeepSeek-V3-0324, Gemini-2.5-Pro, and Qwen3-235B-A22B—form a tightly clustered middle tier with comparable scores, ranging from $47.90$ to $51.32$. Qwen3-30B-A3B again lags behind, a result attributable to its limited model size.
@@
The Impact of an Expanded Action Space. As shown in Figure 4, when transitioning from the Oracle mode to the Standard mode, model performance can be categorized into four distinct groups. (1) Performance Increase: Only one model, Claude-4-Sonnet, benefits from the enlarged action space. Its average score rises from $57.77$ to $61.01$, a relative increase of $5.61\%$. The performance gains are concentrated on more difficult tasks, with improvements of $5.80\%$ on L2 and $9.38\%$ on L3, compared to a modest $2.40\%$ on L1 tasks. (2) Stable Performance: Gemini-2.5-Pro and DeepSeek-R1-0528, belong to this category, showing minimal performance changes. Their scores dropped by a negligible $0.92$ and $0.49$ points, respectively, indicating robustness to the expanded toolset. (3) Moderate Performance Drop: This group includes Qwen3-235B-A22B, Qwen3-30B-A3B, and DeepSeek-V3-0324, which exhibit a more pronounced erosion in accuracy. Their scores declined by $5.64$, $10.53$, and $11.97$ points. We attribute these declines to the increased difficulty of tool selection over a larger set and the substantially heavier context loads compared to the Oracle setting. (4) Significant Performance Drop: GPT-4o-20241120 and Kimi-K2-0711 incur the largest performance drops, with their scores falling by $24.49$ and $39.50$ points. These two models use prompt-based function calling in Standard mode, which may have a negative impact on performance and will be discussed in the following subsection.
@@
Figure 5: A case study on the impact of an expansive action space on Claude-4-Sonnet. Blocked by a strict authentication format in the limited Oracle mode, the model succeeds in Standard mode by pivoting to an alternative tool textttfetch from the expanded toolset.
@@
The performance increase for Claude-4-Sonnet suggests that, given sufficient context-handling capacity, the agentic model can absorb richer tool descriptions and, when blocked, actively pivot to alternative tools and strategies instead of stalling. Figure  5 presents a case study that illustrates how Claude-4-Sonnet turns an expanded toolset from a potential liability into a clear advantage. The figure shows three distinct solution pathways for the user’s task. The annotated ”correct” path (green) involves using the validate and filter tools. In the constrained Oracle mode (red path), Claude-4-Sonnet attempts this path but fails. It provides an incorrect parameter type for auth_token—passing a string instead of the expected boolean—which results in an authorization error and causes the model to abandon the task. However, in the Standard mode (blue path), Claude-4-Sonnet demonstrates adaptive problem-solving. After encountering the same initial failure, it leverages the larger available toolset to pivot to an entirely different tool, fetch. By using this alternative approach, Claude-4-Sonnet successfully resolves the problem.
The performance gap between the Oracle and Standard modes validates the effectiveness of MCPVerse. While most models perform adequately in the constrained Oracle mode, the Standard mode serves as a critical filter. It effectively distinguishes models with genuine agentic tool-use capabilities from those that falter when confronted with a larger and more complex action space.
Prompt-Based Function Calling vs. Native Function Calling. To bypass native tool number limitations, we employed a prompt-based function calling method that integrates tool schemas directly into the system prompt. A comparison (Table 3) of this prompt-based method (Prompt) with native function calling (FC) reveals two key findings.
First, Claude-4-Sonnet exhibited a significant performance degradation in both Oracle and Standard modes. We attribute this to a substantial mismatch between the function-calling templates in our prompt and those from the model’s original training. This discrepancy caused a high rate of hallucination, exceeding 70%, where the model frequently fabricated tool responses despite explicit instructions to the contrary (see Appendix 3 for details).
@@
Claude-4-Sonnet $57.81$ $18.33$ $\downarrow$ $61.01$ $16.65$ $\downarrow$
DeepSeek-V3-0324 $49.55$ $50.07$ $\uparrow$ $37.58$ $30.79$ $\downarrow$
Qwen3-235B-A22B $47.90$ $46.99$ $\downarrow$ $42.26$ $30.61$ $\downarrow$
Qwen3-30B-A3B $33.31$ $35.75$ $\uparrow$ $22.78$ $22.32$ $\downarrow$
GPT-4o-20241120 $51.32$ $45.95$ $\downarrow$ {-} {-}
Table 3: Performance comparison between native function calling (FC) and a prompt-based method (Prompt). The arrows indicate the performance change of the prompt-based method relative to native FC, with green signifying an improvement and red a decline.
Second, for all other models, the performance difference between the Prompt and FC methods was marginal in the constrained Oracle mode. This suggests that models can effectively execute function calls via system prompts when the number of tools is small. In the more complex Standard mode, however, these same models experienced a sharp performance decline. This indicates that as context length increases, their instruction-following capabilities weaken, and they struggle to resolve conflicts between our external formatting instructions and their internal, pre-trained templates.
Model Oracle Mode Standard Mode
L1 L2 L3 L1 L2 L3
Claude-4-Sonnet $\mathbf{7.3}$ $9.8$ $13.6$ $\mathbf{8.2}$ $\mathbf{12.3}$ $\mathbf{14.3}$
DeepSeek-V3-0324 $6.2$ $9.8$ $9.8$ $6.0$ $8.3$ $9.3$
DeepSeek-R1-0528 $6.2$ $8.3$ $11.9$ $7.2$ $9.7$ $11.8$
GPT-4o-20241120 $6.8$ $\mathbf{10.5}$ $12.2$ $5.3$ $3.8$ $4.2$
Qwen3-235B-A22B $6.0$ $8.5$ $10.7$ $5.7$ $9.0$ $8.9$
Qwen3-30B-A3B $6.3$ $8.6$ $10.8$ $6.7$ $8.7$ $7.8$
Gemini-2.5-Pro $5.2$ $6.7$ $8.3$ $5.8$ $7.1$ $7.6$
Kimi-K2-0711 $5.6$ $9.8$ $\mathbf{14.0}$ $3.8$ $4.6$ $3.2$
Table 4: Dialogue length distribution of model responses for Oracle and Standard modes.
Dialogue Length Analysis
Table 4 demonstrates the average number of dialugue length among different models across varying task complexities and tool scale. The data indicate that dialogue length scales with task complexity in both settings. In Standard mode, the mean dialogue length grows from $6.1$ (L1) to $7.9$ (L2) and $8.4$ (L3); in Oracle mode it rises from $6.2$ to $9.0$ and $11.4$, respectively.
A dialogue ends either with a solution or with abandonment after repeated failed tool calls. In Standard mode, Claude-4-Sonnet shows the longest exchanges especially on harder problems, surpassing the next-best model’s dialogue length is reduced by $2.6$ on L2 and $2.5$ on L3. These extra turns indicate a more systematic search, likely boosting its success rate, whereas other models terminate earlier on complex tasks.
Conclusion

### v2

@@ 3.3 Evaluation Modes
  • Standard Mode: This mode is designed for a 64k-token context length, a standard supported by most state-of-the-art LLMs. The toolset in this mode is the union of the minimal tool sets required for each task in Oracle Mode. It consists of $32$ MCPs (totaling $220+$ tools), whose combined definitions occupy approximately 44k tokens, leaving about 20k tokens for the user prompt and the model’s response.
@@
  • Max-Scale Mode: This mode expands on the Standard Mode by loading all $65$ MCPs with $550+$ tools simultaneously. The combined context length required is approximately 140k tokens. In our evaluation, only models with extended context capabilities, such as Claude-4-Sonnet and Gemini-2.5-Pro, were able to complete the evaluation successfully, as they can handle the large number of tools and the extensive context length.
3.4 Evaluation Pipeline
@@
We adopt a simple evaluation pipeline(see Appendix B) with the goal of assessing the LLM’s fundamental capabilities, rather than evaluating the effectiveness of complex agentic frameworks like ReAct (Yao et al., 2023) or RolePlaying (Hu et al., 2025). These advanced methods often depend on sophisticated prompt engineering, which can obscure the model’s intrinsic abilities. In contrast, our direct approach reduces the impact of such artifacts, yielding a more faithful measure of the model’s core agentic capability.
3.5 Task Formulation
Input and Output Consider a collection of MCP servers $S=\{s_{1},s_{2},\ldots,s_{k}\}$, each server $s_{i}$ providing a toolset $T_{i}=\{t_{i,1},t_{i,2},\ldots,t_{i,N}\}$ accessible via the MCP protocol. Given a task $Q$ and a set of candidate data files $D$, the task in MCPVerse is to enable an agent $\mathcal{G}$ to address $Q$ by invoking suitable tools from the available pool $T$, which yields an outcome $\mathcal{O}$, i.e., $\mathcal{O}=\mathcal{G}(Q,D\mid T)$, where $\mathcal{O}$ may represent either a final answer or modifications to the associated file states.
Evaluation Metric Given MCPVerse’s expansive action space, multiple valid action trajectories may exist to accomplish a task. While we cannot directly evaluate every possible trajectory, the final outcome (or end state) must align with the ground-truth specification. Formally, for a task $\tau$ with ground-truth outcome $O$, let $\hat{O}=\mathcal{G}(Q,D\mid T)$ denote the outcome produced by an agent $\mathcal{G}$. We define an evaluation function
$E(\hat{O},O)\in\{0,1\},$
which returns $1$ if the predicted outcome $\hat{O}$ is consistent with the ground truth $O$, and 0 otherwise.
In practice, $E(\cdot)$ is instantiated differently depending on the task type: (i) for tasks with textual ground truth, we employ an LLM and prompt to assess semantic equivalence; (ii) for tasks involving file modifications or state changes, we use a dedicated evaluation script. More details can be found in the Appendix D If $E(\cdot)$ returns $1$, the task is considered successfully completed. Aggregating success rates across all tasks yields a quantitative measure of the model’s agentic capability.
4 EXPERIMENT
4.1 Setups
Our evaluation system, built upon the CAMEL framework (Li et al., 2023a) (Figure 1), conducts evaluations of leading LLMs, including DeepSeek-V3, DeepSeek-V3.1-Terminus, DeepSeek-R1-0528, Claude-4-Sonnet, Qwen3-235B-A22B, Qwen3-235B-2507, Qwen3-30B-A3B, Gemini-2.5-Pro, GPT-4o-20241120 and GPT-5, Kimi-K2-0711 and GLM-4.5 across three distinct modes, as described in subsection 3.3. Due to context length limitations, the Max-Scale mode, which requires approximately 147k tokens, can only be evaluated with models such as Claude-4-Sonnet and Gemini-2.5-Pro. Additionally, for scenarios where the toolset exceeds the API’s tool number limits but does not exceed the context length—such as GPT-5, which caps the number of tools at 128128—we adopt an approach similar to Patil et al. (2025), referred to as prompt-based function calling, embedding the complete tool definitions within a custom system prompt (provided in Appendix  E), thus bypassing the API’s fixed tool number limitations. More setup details see Appendix C.
Table 2: Success rates (SR) of all models in Oracle, Standard, and Max-Scale modes, with breakdowns by task levels (L1–L3). SR is the average of L1, L2, and L3. Context Window denotes the maximum input length (in tokens), and Max Tool Count the maximum number of tools an API allows. In the Tool Count column, a dash (-) means either no tool cap or a limit larger than the full Max-Scale toolset. In the Max-Scale results, a dash indicates the model could not be evaluated due to context constraints. Scores with were obtained via prompt-based function calling.
Model
Context
Window
Max Tool
Count
Oracle Mode Standard Mode Max-Scale Mode
SR L1 L2 L3 SR L1 L2 L3 SR L1 L2 L3
DeepSeek-V3-0324 $64\mathrm{k}$ $128$ $46.8$ $62.7$ $45.1$ $32.7$ $32.2$ $47.0$ $27.7$ $22.0$ {-} {-} {-} {-}
DeepSeek-V3.1-Terminus $\mathrm{k} 128128$ $56.6$ $64.4$ $57.3$ $48.3$ $52.1$ $62.0$ $49.4$ $45.0$ {-} {-} {-} {-}
DeepSeek-R1-0528 $64\mathrm{k}$ $128$ $56.4$ $70.5$ $56.4$ $42.4$ $49.9$ $65.1$ $47.4$ $37.3$ {-} {-} {-} {-}
Claude-4-Sonnet $200\mathrm{k}$ - $62.3$ $71.6$ $62.7$ $52.5$ $\mathbf{62.4}$ $\mathbf{75.9}$ $60.4$ $\mathbf{50.9}$ $\mathbf{44.2}$ $\mathbf{45.8}$ $\mathbf{40.5}$ $\mathbf{46.2}$
Qwen3-235B-A22B $128\mathrm{k}$ - $42.1$ $61.6$ $34.8$ $30.0$ $37.7$ $52.3$ $35.9$ $25.0$ {-} {-} {-} {-}
Qwen3-235B-2507 $1\mathrm{M}$ - $44.8$ $62.5$ $44.8$ $27.1$ $53.2$ $63.9$ $51.8$ $43.9$ $31.6$ $44.3$ $23.7$ $26.7$
Qwen3-30B-A3B $128\mathrm{k}$ - $27.7$ $46.5$ $18.1$ $18.3$ $18.9$ $27.9$ $12.9$ $15.8$ {-} {-} {-} {-}
Gemini-2.5-Pro $1\mathrm{M}$ $512$ $48.7$ $66.3$ $42.6$ $37.3$ $45.3$ $62.4$ $41.9$ $31.6$ $31.4^{*}$ $37.7^{*}$ $35.7^{*}$ $20.8^{*}$
GPT-4o-20241120 $\mathrm{k} 128128$ $42.1$ $59.1$ $40.2$ $27.1$ $31.4^{*}$ $37.7^{*}$ $35.7^{*}$ $20.8^{*}$ {-} {-} {-} {-}
GPT-5 $400\mathrm{k}$ - $\mathbf{68.1}$ $\mathbf{80.5}$ $\mathbf{67.8}$ $\mathbf{55.9}$ $23.4^{*}$ $30.6^{*}$ $19.7^{*}$ $20.0^{*}$ {-} {-} {-} {-}
Kimi-K2-0711 $\mathrm{k} 128128$ $59.4$ $70.9$ $59.8$ $47.5$ $16.3^{*}$ $^{*} 24.424.4^{*}$ $0.0^{*}$ {-} {-} {-} {-}
GLM-4.5 $128\mathrm{k}$ - $55.0$ $70.9$ $58.5$ $35.6$ $59.1$ $67.4$ $\mathbf{60.7}$ $49.2$ {-} {-} {-} {-}
4.2 Results Analysis
@@
This section presents a detailed analysis of the experimental results. Owing to tool-mounting constraints, the Standard mode score for GPT-4o-20241120 and GPT-5 and the Max-Scale mode score for Gemini-2.5-Pro are both obtained via prompt-based function calling.
4.2.1 Overall Model Performance
(1) Winners by mode. From Table 2, the best Oracle SR is achieved by GPT-5 (68.1), whereas Standard and Max-Scale are led by Claude-4-Sonnet with 62.4 and 44.2, respectively. GLM-4.5 achieve a very close performance to Claude-4-Sonnet.
(2) MCPVerse is Challenging. On the one hand, as shown in Table 2, even the strongest models achieve modest scores—under Max-scale mode the top model (Claude-4-Sonnet) reaches only 44.2 average accuracy, and on the most difficult L3 subset the best score is just 46.2. On the other hand, many models cannot be run end-to-end under normal settings: Max-Scale demands $\sim$147k tokens and is therefore feasible only for Claude-4-Sonnet and Gemini-2.5-Pro; moreover, several APIs cap the number of tools (e.g., 128 for GPT-4o-20241120/GPT-5 and 512 for Gemini), preventing the full Standard/Max-Scale toolsets from being mounted natively.
4.2.2 The Impact of an Expanded Action Space
@@
As shown in Fig. 3, models exhibit heterogeneous trends when moving from Oracle to Standard: three improve, whereas the remainder decline, with drops ranging from mild to severe. (1) Improvements trace recent agentic advances. The three gainers—Claude-4-Sonnet, Qwen3-235B-2507, and GLM-4.5—are recent releases within the past four months, positioned for agentic use; their higher Standard scores indicate stronger exploration and exploitation of a large-scale tool space, showing that MCPVerse sensitively captures progress in agentic capability. (2) Severe declines align with API tool caps. Models with large drops (GPT-4o-20241120/GPT-5/Kimi-K2-0711)—highlighted by dashed traces—share a common constraint: API tool-number limits prevent native function calling in Standard mode, forcing the use of prompt-based function calling. This alternative mechanism likely accounts for their performance degradation and will be discussed in detail in Section 4.2.4. (3) Mild declines reflect context burden. For the remaining models not in (1) or (2), performance degrades as the toolset grows, consistent with heavier schema context and harder tool selection in the larger action space. (4) Larger action spaces remain challenging. For those evaluable at Max-Scale, performance further decreases from Standard to Max-Scale, indicating that even top systems still struggle as the tool space expands.
4.2.3 Emergent Solution Paths and “Hacking” Behaviors
@@
Figure 4: Case study: Claude-4-Sonnet discovers a new solution path. Blocked by a strict authentication format in the limited Oracle mode, the model succeeds in Standard mode by pivoting to an alternative tool textttfetch from the expanded toolset.
@@
Performance gains from Oracle to Standard for several models point to emergent solution paths in larger action spaces: with more tools available and sufficient context budget, models are better able to explore, compose, and exploit alternative tool chains rather than stall on a blocked route. Figure 4 presents a case study in which Claude-4-Sonnet turns an expanded toolset from a potential liability into a clear advantage. The figure depicts three distinct pathways for the same user task. The annotated path (green) is the canonical route, using validate and filter from the dataviewer MCP server. In the constrained Oracle setting (red), Claude-4-Sonnet initially follows this route but fails due to a type mismatch on auth_token—it supplies a string instead of the required boolean—triggering an authorization error and causing the attempt to be abandoned. In the Standard setting (blue), after encountering the same initial failure, the model leverages the larger toolset to pivot to fetch and completes the task. This switch illustrates the ‘hacking’ behavior we observe: the model reaches the goal by switching to another solution path and bypassing the interface that had blocked the annotated path.
4.2.4 Prompt-Based Function Calling vs. Native Function Calling
Table 3: Performance comparison between native function calling (FC) and a prompt-based method (Prompt)
@@
Claude-4-Sonnet $62.28$ $35.55$ $\downarrow$ $62.36$ $15.10$ $\downarrow$
DeepSeek-V3-0324 $46.80$ $45.79$ $\downarrow$ $32.23$ $24.92$ $\downarrow$
DeepSeek-V3.1-Terminus $56.64$ $55.05$ $\downarrow$ $52.13$ $41.43$ $\downarrow$
Qwen3-235B-2507 $44.82$ $49.66$ $\uparrow$ $53.18$ $37.99$ $\downarrow$
Qwen3-235B-A22B $42.12$ $45.49$ $\downarrow$ $37.74$ $26.94$ $\downarrow$
Qwen3-30B-A3B $27.65$ $28.86$ $\uparrow$ $18.88$ $20.67$ $\uparrow$
GPT-4o-20241120 $42.13$ $42.43$ $\uparrow$ {-} {-}
GPT-5 $68.06$ $68.28$ $\uparrow$ {-} {-}
Kimi-K2-0711 $59.41$ $60.57$ $\uparrow$ {-} {-}
GLM-4.5 $55.00$ $52.68$ $\downarrow$ {-} {-}
To bypass native tool number limitations, we employed a prompt-based function calling method that integrates tool schemas directly into the system prompt. A comparison (Table 3) of this prompt-based method (Prompt) with native function calling (FC) reveals two key findings. First, Claude-4-Sonnet exhibited a significant performance degradation in both Oracle and Standard modes. We attribute this to a substantial mismatch between the function-calling templates in our prompt and those from the model’s original training. This discrepancy caused a high rate of hallucination, exceeding 70%, where the model frequently fabricated tool responses despite explicit instructions to the contrary. Second, for all other models, the performance difference between the Prompt and FC methods was marginal in the constrained Oracle mode. This suggests that models can effectively execute function calls via system prompts when the number of tools is small. In the more complex Standard mode, however, these same models experienced a sharp performance decline.
4.3 Evaluation under Retrieve Mode
In the MCPVerse evaluation system, although the tool space is already large, we initially excluded retrieval because we believed it could restrict the LLM’s exploration of the full tool space. In this section, we add experiments under the retrieve mode. Following the semantic similarity-based retrieval methods of Patil et al. (2024) and Moon et al. (2024), we use OpenAI’s text-embedding-3-large model to embed the descriptions of all MCP servers and store them in a database. During evaluation, each tested model first generates the descriptions of the MCP servers it needs. These descriptions are also embedded with text-embedding-3-large, and the top-5 most relevant servers are selected using cosine similarity. These retrieved servers are then mounted to the model for MCPVerse evaluation. The key difference between retrieve mode and oracle mode lies in the toolset that is mounted. In oracle mode, the annotated ground-truth servers are provided, whereas in retrieve mode, more servers may be mounted but the selection is less precise. Under this setting, as shown in Table 4, model performance in retrieve mode is consistently and significantly lower than in oracle mode and standard mode.
Table 4: Comparison of model success rates (SR) across Oracle, Standard, and Retrieve modes.
Model Oracle (SR) Standard (SR) Retrieve (SR)
DeepSeek-V3.1-Terminus 56.6 52.1 39.8
Qwen3-235B-2507 44.8 53.2 39.4
Claude-4-Sonnet 44.8 53.2 40.3
GPT-5 68.1 23.4 45.0
5 Conclusion